I wanted to make sure that I have the correct answer, can you check this one please.
solve the following system of equations
x+4y=8 (1)
x=4-4y (2)
What is the solution of the system?
I put the solution is (1,2)ordered pair. Is this correct?

Answer by Alan3354(69443)

I wanted to make sure that I have the correct answer, can you check this one please.
solve the following system of equations
x+4y=8 (1)
x=4-4y (2)
What is the solution of the system?
I put the solution is (1,2)ordered pair. Is this correct?
--------------
To check, sub 1 & 2 for x & y.
It doesn't satisfy either equation.
x+4y=8 (1)
x=4-4y (2) --> x+4y = 4
These are inconsistent. If you graph them, you get parallel lines.
They do not intersect, so there is no solution.